Please help, test tomorrow
(2x+5)^1/2=4

Answer by stanbon(48564)

(2x+5)^1/2=4
(2x+5)^1/2=4
-----------
Square both sides to get:
2x+5 = 16
2x = 11
x = 11/2
==============
Cheers,
Stan H.